OpenStudy (anonymous):

simplify: sqrt6 (7sqrt3 + 6)

OpenStudy (anonymous):

@CrossFaded

OpenStudy (jdoe0001):

\(\Large \bf \sqrt[6]{7\sqrt[3]{3+6}} \quad ?\)

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

what does the s stand for next to the 7?

OpenStudy (jdoe0001):

actually, I meant... \(\Large \bf \sqrt[6]{7\sqrt[3]{6}}\quad ?\)

OpenStudy (jdoe0001):

s?

OpenStudy (anonymous):

no the 3+6 are in parantheses than they have a sqrt symbol....hold on I will attach a pic

OpenStudy (jdoe0001):

\(\Large \bf \sqrt[6]{7\sqrt{3+6}}\quad ?\)

OpenStudy (jdoe0001):

\(\bf \sqrt{6}(7\sqrt{3}+6)\qquad distribute\\ \quad \\ (\sqrt{6}\cdot 7) + (\sqrt{6}\cdot 6)\)

OpenStudy (jdoe0001):

wooops,, my 3 went awol =) hehehe \(\bf \sqrt{6}(7\sqrt{3}+6)\qquad distribute\\ \quad \\ (\sqrt{6}\cdot 7\sqrt{3}) + (\sqrt{6}\cdot 6)\)

OpenStudy (anonymous):

ok thank you

OpenStudy (jdoe0001):

\(\bf \sqrt{6}(7\sqrt{3}+6)\qquad distribute\\ \quad \\ (\sqrt{6}\cdot 7\sqrt{3}) + (\sqrt{6}\cdot 6)\implies (7\sqrt{6\cdot 3})+6\sqrt{6} \)